# Improving Human-AI Coordination through Online Adversarial Training and Generative Models

### v3

@@ 4 GOAT: Generative Online Adversarial Training
Cross-Play (XP): We begin by introducing an adversarial training objective which we term cross-play, following its use in other work in PBT-based cooperation methods [4]. Here, an Adversary policy $\pi_{A}$ attempts to minimize cooperation performance, while a Cooperator policy $\pi_{C}$ attempts to maximize it. This type of optimization is widely known as minimax optimization in game theory:
@@
We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation 1. Now, the Adversary policy conditions on a randomly sampled $z\sim\mathcal{N}(0,I)$, and transforms it to produce an Adversary embedding $z^{\prime}=\pi_{A}(z)$. This modified $z^{\prime}$ is fed to the decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$ of the generative model to produce the actions of the simulated partner policy $\pi_{P}$. Thus, the optimization problem is as follows:
@@
$=$ $\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}[V(\pi_{C},\pi_{P})]$ (3)
However, this minimax objective still does not guarantee effective exploration in the partner space as it optimizes for worst-case returns. Therefore, the Adversary will only search for the worst agents that the generative model can simulate. Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies which could have good performance, but the Cooperator failed to cooperate well with them.
Regret-based adversarial cooperative training. To address the shortcomings of the minimax objective, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent. We now formalize the definition of Regret we will use for cooperative settings.
Self-Play (SP): This play style, first used by Samuel [27] to play checkers, is defined as the expected return when an agent is paired with itself:
$J_{SP}=\mathbb{E}[V(\pi_{P},\pi_{P})]$ (4)
Regret ($R$): As proposed initially by Robbins [25], regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight. We formalize regret in the cooperative setting as the performance gap between the cross-play (XP) performance an agent achieves with a partner $\pi_{P}$, and $\pi_{P}$’s optimal self-play (SP) performance.
@@
At the start of the training process, the GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$. This latent vector $z^{\prime}$ is then used to generate the partner policy $\pi_{P}$ using the frozen weights of the VAE decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$. The partner policy $\pi_{P}$ is then paired with itself to evaluate SP returns and with $\pi_{C}$ to evaluate XP returns. Regret is evaluated after collecting the returns from both games. The Adversary is trained to choose $z^{\prime}$ to maximize regret, while the Cooperator policy is trained to maximize its cooperation performance. The full objective of the GOAT game is thus:
$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P},\pi_{C})$ (6)
The online optimization as shown in algorithm 1 ensures that the Adversary is incentivized to continually discover new partner strategies that are challenging for the Cooperator, by searching the latent space to maximize regret. The Cooperator in turn learns to adapt to these partner strategies by minimizing the regret. This cyclic behavior naturally trains the Cooperator to learn to cooperate with diverse strategies, from more straightforward to increasingly complex ones. For invalid partners, the optimal policy cannot achieve positive returns, resulting in zero regret and providing no incentive to the Adversary. This property of regret promotes exploration that leads to curriculum generation to increasingly generate complex yet solvable tasks to train the Cooperator agent.
@@ 5 Experiments
In this section, we describe our experimental evaluation, which will allow us to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans. We evaluate GOAT using two popular coordination benchmarks: 1)1)1 ) One-Step Cooperative Matrix Game (CMG) [20, 43, 4] and 2)2)2 ) Overcooked [3].
Figure 2: Carroll et al. [3] first introduced the first five layouts: Cramped Room, Asymmetric Advantages, Coordination Ring, Forced Coordination, and Counter Circuit. To increase coordination complexity, Liang et al. [18] introduced the Multi-Strategy Counter layout, where players could choose between tomato and onion ingredients to cook soup. This adds an additional coordination challenge where agents have to adapt to soup-making strategies based on different ingredients.
5.1 One-Step Cooperative Matrix Game (CMG):
@@
This is a single-step game where two agents (Row agent and Column agent) independently select an action by choosing a row and a column, respectively, and the intersection of their choices returns the associated reward according to the payoff matrix. Both agents receive the same reward and attempt to maximize the common payoff. A typical payoff matrix in CMG is defined by $(M,\{k_{m}\},\{r_{m}\})$ where $M$ is number of solutions, $k_{m}$ is number of compatible actions, and $r_{m}$ is the reward associated with each solution. We use a challenging multimodal layout called CMG-M as shown in Figure 3.
5.2 Overcooked:
@@
This is a collaborative game where two players need to collaborate by dividing cooking tasks and avoiding obstructing each other. Identifying the opposite player’s intent and behavior is key to determining how to coordinate with the other agent effectively. We test GOAT on the first five layouts introduced in [3], and add a more complex layout introduced by Liang et al. [18], the Multi-Strategy Counter (all shown in Figure 2).
5.3 GOAT implementation & Baselines:
@@
The Cooperator is trained using PPO, and the Adversary is trained using REINFORCE. Training the Adversary is RL-algorithm-agnostic as it does not depend on state and is a one-step optimization process. Further details on the KL-regularization of the VAE, training procedure for GOAT, and other architecture and hyperparameters are available in the Appendix. We compare to 5 competitive prior baselines on human-AI coordination: Behavior Cloning (BC) partners with RL cooperator [3] (BC+RL), PBT methods Fictitious Co-Play (FCP) [32], maximum entropy population (MEP) [43], CoMeDi [28], and GAMMA trained on MEP (MEP+GAMMA) [18], which previously showed state-of-the-art performance in real human evaluations.
5.4 Human Evaluation:
@@
We conducted a user study with 40 participants to discover which method best coordinated with humans and is evaluated as qualitatively better. Because we want the most challenging tasks with the most diverse strategy space, we focus on only the most complex game layouts, Counter Circuit, and Multi-strategy Counter. The latter presents a more complex cooperation task because partners can prepare two dishes using different ingredients and thus represents a more rigorous test of whether algorithms can generalize to cooperate with real humans. Participants were recruited from the Prolific online crowdsourcing platform following an IRB-approved protocol. Each participant was tasked with completing 6 total rounds of the Overcooked game with an AI agent partner, where each round presented a different agent model among one of the baselines or our method. In order to minimize trends due to gameplay order, the agent models are loaded in randomized order, with a randomly sampled checkpoint chosen from 1 of 5 random seeds.
@@ 6 Experimental Results
(a) CMG Layout
(b) Self-play
(c) TrajeDi
(d) MEP
(e) CoMeDi
(f) GOAT
Figure 3: To test the ability of the method to discover and exploit multiple minima and maxima in the problem space, we introduce a custom CMG layout (a). This reward structure is designed to have three scenarios: no reward, suboptimal rewards, and global maximum reward. The policies for the methods were aggregated and normalized into a heatmap.
6.1 RQ1: Can GOAT discover and exploit multiple cooperative modes?
For GOAT to be effective, the VAE should be able to discover and exploit regions of the reward space. To test this, we use a one-step cooperative matrix game (CMG). We test the cooperative policy generation capability of VAE by training it on trajectory data generated by the population of CoMeDi agents. Since CoMeDi data is cooperative, we hypothesize that the VAE should only generate policies that are in distribution. To keep the evaluation fair, we used a population size of 8 for all the methods tested on CMG-M. During training, GOAT covers all the modes and correctly assigns importance to the agent’s actions. As seen in Figure 3, the baselines self-play, Trajedi, and MEP get stuck and might not discover other modes. CoMeDi discovers all the regions and gives the highest priority to the global maximum, but spreads probability mass roughly equally across all other modes, regardless of their payoff. In contrast, GOAT not only covers all modes but is better at assigning higher probability to more rewarding modes (for example, assigning low probability to modes 4 and 7, which have low payoff).
We note that the quality of VAE training and the amount of training data greatly influence the resulting GOAT policy. To robustly train the VAE, we follow Liang et al. [18] and augment the 8 policy pool data by pairing the row and column agents of one policy with other policies. In our experiments, we saw that higher $\beta$ for KL regularization in VAE resulted in better generation of actions.
6.2 RQ2: Will adversarial training result in a more robust Cooperator with better sample complexity?
@@
PBT methods, which rely on simulated populations of self-play partners, are often computationally expensive and have poor coverage of actual human behaviors. Training a Cooperator against them may not achieve good performance with people, and it may take a long time since the training does not focus on correcting weaknesses in the Cooperator but instead randomly samples from the population. We hypothesize that our adversarial training technique can more efficiently search the space of partners to make an agent that is more robust and requires less training time.
To validate this hypothesis, we investigate the learning curves of all simulated training methods, including our own, in Figure 4. Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work [3, 32, 18]. The human proxy enables the automatic evaluation of agents’ performance throughout training.
@@
As is evident in the curves, GOAT achieves significantly higher performance than all 4 baselines in both layouts and increases performance more quickly than other techniques. We hypothesize this is because the generative model, trained on a broad spectrum of cooperative trajectories, encodes a latent space rich in strategic variations. Adversarial exploration of this space enables rapid coverage of human behaviors ranging from simple coordination patterns to complex, context-dependent strategies. By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies while still capturing the semantic diversity critical for generalizing to diverse partners. Because it efficiently targets training to weak points in the Cooperator’s policy, it has lower sample complexity than traditional PBT methods.
Figure 4: Evaluation of GOAT against a simulated human policy trained with behavior cloning (BC). We compare to 4 baselines which are trained using simulated population data, including the previous state-of-the-art method, GAMMA [18]. Error bars show the standard error over 5 random seeds.
6.3 RQ3: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans?
@@
As described in Section 5, we evaluated the Cooperator agents trained with GOAT and the 5 baselines alongside real human partners in the Counter Circuit and the Multi-strategy Counter layouts. We recruited novel human players whom the agents had not previously seen. Figure 5 depicts the evaluation results for the two layouts and all the agents against human players. On the Counter Circuit layout, we find that the performance of GOAT is not significantly different from the previous state-of-the-art method, GAMMA (see Table 4 for statistical significance results). We hypothesize this is because both methods are reaching close to optimal performance on this layout. However, on the more complex layout, Multi-Strategy Counter, the GOAT leads to significantly higher cooperation performance with real humans, leading to a 38% improvement over the previous state of the art. To test these differences for yourself, please visit our webpage (https://sites.google.com/view/goat-2025/), which provides an interactive demo in which you can play Overcooked with both GOAT and the baselines.
Figure 5: Evaluation of Performance of different methods when tested against real humans in two layouts: counter circuit (left plot) and multi-strategy counter (right plot).
6.4 RQ4: Will adversarial training create a curriculum of increasingly challenging strategies by discovering them in latent space?
The regret-driven adversarial framework is designed to create a dynamic curriculum where the Cooperator must continually adapt to novel challenges. GOAT identifies underperforming coordination strategies and generates increasingly complex cooperation challenges, compelling the Cooperator to explore and master new behaviors. This iterative process prevents the Cooperator from stagnating into rigid conventions, a common pitfall in self-play or population-based methods. Here, we assess how GOAT explores the latent space of the generative model over the whole training interval.
Figure 6: Left: GOAT exploring the latent space exploration compared against GAMMA [18] in red over the whole training period. Right: Minimax Adversary in red and GOAT in blue exploring the latent space over the training period. We provide a visual example related to this on our webpage.
@@
The latent vector embeddings chosen by the GOAT over the whole training interval are shown in Figure 6. We compare this to the distribution of embeddings of the original GAMMA method [18] samples throughout training (red ball). We can see that at the start of training, the GOAT samples latent vectors distributed throughout this ball (shown in purple). As training progresses, it chooses a particular region to exploit. After a few iterations, it moves to a new region, presumably because the Cooperator adapts to that strategy, and it is no longer able to maximize regret. Finally, at the end of the training step in yellow, GOAT has traveled across the latent space to explore particular regions (those with high regret) to find valid strategies that exploit weaknesses in the Cooperator.
6.5 RQ5: Is minimax objective enough for adversarial training, and how well does it perform against the regret objective?
@@
In Section 4, we described how the purely adversarial minimax objective (Equation 1) is misaligned with effectively training agents to cooperate. In Equation 3, we showed how integrating a generative model could help alleviate this problem. However, while the generative model can prevent sabotage behavior, that does not mean all agents it simulates will show productive cooperation behavior. We claimed that the purely adversarial minimax objective would still focus too much on poorly performing or incompetent partners, motivating the use of the regret objective in Equation 5. We now empirically support this claim by comparing GOAT to an ablation of our method, which uses the minimax objective in Equation 3, on the Multi-Strategy Counter layout. As we can see in Figure 7, the GOAT outperforms the minimax trained Cooperator, suggesting that using a regret-based objective is an important component of GOAT. We further investigate the embeddings chosen by both types of Adversary throughout the training period in Figure 6 (right). We find that the minimax Adversary mostly fixates on a single region (red cluster in the plot) where the policies result in low reward behaviors. One such example of the adversarial strategies is where the resulting partner agent stands still in a corner to satisfy its objective to minimize performance, but makes no attempts to cooperate. In the blue, we have embeddings chosen by the GOAT throughout the training period. We can observe that the regret objective causes the GOAT to move to explore multiple modes over the course of training, exploring the latent space more effectively and thus covering a wider range of interesting behaviors.

### v4

@@ 4 GOAT: Generative Online Adversarial Training
Adversarially minimizing Cross-Play (XP) performance. We begin by introducing an adversarial training objective based on the XP performance of the Cooperator $\pi_{C}$ and an adversarial partner $\pi_{A}$. The adversary $\pi_{A}$ attempts to minimize cooperation performance, while a Cooperator policy $\pi_{C}$ attempts to maximize it. This type of optimization is widely known as minimax optimization in game theory:
@@
We formalize this adversarial training algorithm by modifying the previous $J_{XP}$ formulation in Eq. 1. Now, the Adversary policy conditions on a randomly sampled $z\sim\mathcal{N}(0,I)$, and transforms it to produce an Adversary embedding $z^{\prime}=\pi_{A}(z)$. This modified $z^{\prime}$ is fed to the decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$ of the generative model to produce the actions of the simulated partner policy $\pi_{P}$. Let $\pi_{P}^{\pi_{A}(z)}$ be the partner policy produced by having the adversary $\pi_{A}$ choose an embedding which is then simulated with the generative model. Then, the optimization problem is as follows:
@@
$=$ $\max_{\pi_{C}}\min_{\pi_{A}}\mathbb{E}_{z\sim\mathcal{N}(0,I)}[V(\pi_{C},\pi_{P}^{\pi_{A}(z)})]$ (3)
This objective no longer directly trains the partner policy on the adversarial objective, but instead uses the adversary to steer the pre-trained generative model to select partners.
Regret-based adversarial cooperative training. Because Eq. 3 is still fundamentally a minimax objective, it does not guarantee effective exploration in the partner space because it optimizes for worst-case returns. Therefore, the Adversary will only search for the worst agents that the generative model can simulate. Instead, we would like to formulate a better objective that encourages searching for meaningful partner policies that could have good performance, but for which the Cooperator still fails to perform well. Thus, we propose using a regret objective to generate a realistic curriculum for the Cooperator agent. As proposed initially by Robbins (1952), regret is the performance gap between the returns of an agent-chosen strategy and the returns of the optimal strategy that could have been selected with perfect foresight. We formalize regret in the cooperative setting as the performance gap between the cross-play (XP) performance an agent achieves with a partner $\pi_{P}$, and $\pi_{P}$’s optimal self-play (SP) performance, $\mathbb{E}[V(\pi_{P},\pi_{P})]$, which is the score it achieves when paired with itself.
@@
GOAT objective. At the start of the training process, GOAT samples $z$ from the standard normal distribution and transforms it to the effective latent vector $z^{\prime}$. This latent vector $z^{\prime}$ is then used to generate the partner policy $\pi_{P}$ using the frozen weights of the VAE decoder $p(a_{t}|z^{\prime},\tau_{0:t-1};\theta)$. The partner policy $\pi_{P}$ is then paired with itself to evaluate SP returns and with $\pi_{C}$ to evaluate XP returns. Regret is evaluated after collecting the returns from both games. The Adversary is trained to choose $z^{\prime}$ to maximize regret, while the Cooperator policy is trained to maximize its cooperation performance. The full objective of the GOAT game is thus:
$\min_{\pi_{C}}\max_{\pi_{A}}Regret(\pi_{P}^{\pi_{A}(z)},\pi_{C})$ (5)
The online optimization, as shown in algorithm 1, ensures that the Adversary is incentivized to continually discover new partner strategies that are challenging for the Cooperator, by searching the latent space to maximize regret. The Cooperator, in turn, learns to adapt to these partner strategies by minimizing the regret. This cyclic behavior naturally trains the Cooperator to learn to cooperate with diverse strategies, from more straightforward to increasingly complex ones. For invalid partners, the optimal policy cannot achieve positive returns, resulting in zero regret and providing no incentive to the Adversary. This property of regret promotes exploration that leads to curriculum generation to increasingly generate complex yet solvable tasks to train the Cooperator agent.
@@ 5 Experiments
Our experimental evaluation is designed to assess how well our adversarial training procedure can improve zero-shot cooperation performance, both in simulation and with real humans.
Environments. We evaluate GOAT using three popular coordination benchmarks: $1)$ One-Step Cooperative Matrix Game (CMG) (Lupu et al., 2021; Zhao et al., 2022; Charakorn et al., 2023), $2)$ Cooperative Reaching Game (CRG) (Rahman et al., 2023; 2024; Charakorn et al., 2023), and $3)$ Overcooked (Carroll et al., 2020).
(a) CMG Payoff Matrix
(b) Self-Play
(c) MEP
(d) CoMeDi
(e) GAMMA
(f) MiniMax
(g) GOAT
(h) Expected Reward
Figure 2: $a)$ Cooperative Matrix Game. $b)\text{ to }g)$ Policy probability distribution to show coverage of different methods on the CMG payoff matrix $h)$ Total expected rewards for each method, assuming we uniformly sample partners and each method gets a payoff for each partner proportional to the amount of coverage they have for the reward block.
@@
  • One-Step Cooperative Matrix Game (CMG). CMG is a single-step game where two agents (Row and Column) independently select an action by choosing a row and a column, respectively, and the intersection of their choices returns the associated reward according to the payoff matrix. Both agents receive the same reward. A typical payoff matrix in CMG is defined by $(M,\{k_{m}\},\{r_{m}\})$ where $M$ is the number of solutions, $k_{m}$ is the number of compatible actions, and $r_{m}$ is the reward associated with each solution. We use a challenging layout shown in figure 2(a) with three scenarios: no reward, suboptimal rewards, and global maximum reward.
  • Cooperative Reaching Game (CRG). CRG is a $5\text{x}5$ grid world, with $|\mathcal{A}|=5$, where two agents need to reach the same goal location simultaneously to receive rewards. For agents to cooperate and receive rewards, they must reach and stay in any of the same goal coordinates. An ideal Cooperator should chase their partner to the same goal or persuade them to follow the same goal. We evaluate GOAT and prior baselines against $11$ heuristic agent teammates described in detail in appendix E.
@@
  • Overcooked. A state-of-the-art collaborative, real-time benchmark where two players need to cooperate by dividing cooking tasks and avoiding obstructing each other. Identifying the opposite player’s intent and behavior is key to determining how to coordinate with the other agent effectively. We test GOAT on the most challenging tasks with a diverse strategy space. Counter Circuit introduced in Carroll et al. (2020), and a more complex layout introduced by Liang et al. (2024), Multi-Strategy Counter.
@@
GOAT implementation & Baselines: Training the Adversary is RL-algorithm-agnostic as it does not depend on state and is a one-step optimization process. To keep the optimization loop simpler, the Cooperator is trained using PPO, and the Adversary is trained using REINFORCE. Details on the KL-regularization of the VAE, training procedure for GOAT, and other architecture and hyperparameters are available in the Appendix. We compare to 5 competitive prior baselines on human-AI coordination: Behavior Cloning (BC) partners with RL cooperator (Carroll et al., 2020) (BC+RL), PBT methods Fictitious Co-Play (FCP) (Strouse et al., 2022), maximum entropy population (MEP) (Zhao et al., 2022), CoMeDi Sarkar et al. (2023), and GAMMA trained on MEP (MEP+GAMMA) Liang et al. (2024), which previously showed state-of-the-art performance in real human evaluations.
@@
Human Evaluation: We conducted a user study with 40 participants to discover which method best coordinated with humans in a real-time evaluation in Overcooked, recording the team scores. Participants were recruited from the Prolific online crowdsourcing platform following an IRB-approved protocol. Each participant was tasked with completing 6 total rounds of the Overcooked game with an AI agent partner, where each round presented a different agent model among one of the baselines or our method. In order to minimize trends due to gameplay order, the agent models are loaded in randomized order, with a randomly sampled checkpoint chosen from 1 of 5 random seeds.
@@ 6 Experimental Results
Figure 3: CRG: average reward obtained against the 11 Heuristic Agent teammates across 5 seeds of each method.
RQ1: Can GOAT discover and exploit multiple cooperative strategies? For GOAT to be effective at generalizing to novel partners, the adversary should be able to discover and exploit all rewarding cooperative modes. We first test this question with the CMG environment, using a population size of 8 for all baseline PBT methods. To train the VAE using the population of CoMeDi agents, we use the same procedure as GAMMA Liang et al. (2024), which pairs each of the 8 row and column agents in the population together to generate trajectory data. Since VAE does not have a population, it can generate many more agents than PBT methods.
As seen in Figure 2(b) and 2(c), the baselines’ Self-Play and MEP fail to cover all modes, implying they would not generalize to partners that used those modes. CoMedi 2(d) and GAMMA 2(e) discover all the regions and give the highest priority to the global maximum, but spread probability mass roughly equally across all other modes, regardless of their payoff. Whereas MiniMax 2(f) converges to the worst-case scenario. In contrast, GOAT not only covers all the modes but also is better at assigning higher probability to more rewarding modes (for example, assigning low probability to modes 3, 4, and 7, which have low payoff). Thus, GOAT achieves good coverage while also focusing on maximizing return. This is also evident in Figure 3, which shows the results of evaluating each method in the CRG environment. Here we see that GOAT achieves the highest average score when teamed with all 11 heuristic partner agents.
(a) Counter Circuit
(b) Simulation Results
(c) Human evaluation
(d) Multi-Strategy Counter
(e) Simulation Results
(f) Human evaluation
Figure 4: Overcooked: Carroll et al. (2020) first introduced the most challenging Counter Circuit layout $(a)$. To increase coordination complexity, Liang et al. (2024) introduced the Multi-Strategy Counter layout $(d)$, where players could choose between tomato and onion ingredients to cook soup. This adds an additional coordination challenge where agents have to adapt to soup-making strategies based on different ingredients. $(b)\text{ \& }(e)$ are evaluation of GOAT against a human proxy model trained with BC. We compare to 4 baselines, which are trained using simulated population data, including the previous state-of-the-art method, GAMMA (Liang et al., 2024). Error bars show the std. err. over 5 random seeds. $(c)\text{ \& }(f)$ shows the evaluation of the performance of methods when tested against real humans in two layouts: counter circuit and multi-strategy counter, respectively.
@@
RQ2: Will adversarial training result in a more robust Cooperator with better sample complexity in simulation? PBT methods rely on simulated populations of self-play partners, are often computationally expensive, and have poor coverage of actual human behaviors. Training a Cooperator against them may not achieve satisfactory performance with people, and it may take a long time to converge since the training does not focus on correcting weaknesses in the Cooperator, but instead randomly samples from the population. We hypothesize that our adversarial training technique can more efficiently search the space of partners to make an agent that is more robust and requires less training time. To validate this hypothesis, we investigate the Overcooked learning curves of all simulated training methods, including our own, in Figure 4(b) and 4(e). Here, we evaluate the cooperation performance of the agents against a Human Proxy model (behavior-cloned (BC) model trained on human data) as in prior work (Carroll et al., 2020; Strouse et al., 2022; Liang et al., 2024). The human proxy enables the automatic evaluation of agents’ performance throughout training.
@@
As is evident in the curves, GOAT achieves significantly higher performance than all 4 baselines in both Overcooked layouts and increases performance more quickly than other techniques. We hypothesize this is because the generative model, trained on cooperative trajectories, encodes a latent space rich in strategic variations. Adversarial exploration in this latent space enables rapid coverage of diverse human behaviors ranging from simple coordination strategies to complex, context-dependent strategies. By constraining the Adversary to valid cooperative policies, the method avoids degenerate strategies and exposes the Cooperator to all the semantically diverse strategies that are beneficial for generalizing to diverse partners. It efficiently targets training to weak points in the Cooperator’s policy, and thus has lower sample complexity than traditional PBT methods.
@@
RQ3: Can GOAT achieve improved cooperation performance when evaluated in real-time with novel humans? The real test of a human-AI coordination algorithm is whether it can work with actual people. As described in Section 5, we evaluated the Cooperator agents trained with GOAT and the 5 baselines alongside real human partners in the Counter Circuit and the Multi-Strategy Counter Overcooked layouts. We recruited novel human players with no prior exposure to the agents. Figure 4(c) and 4(f) depict the evaluation results for the two layouts and all the agents against human players. GOAT outperforms the previous baselines on both layouts. Even though GOAT shows significant improvement over baselines on Counter Circuit simulation results, the magnitude of the improvement over GAMMA is smaller ($3\%$) on this layout. This is because both methods are reaching close to optimal performance in this simpler environment. However, on the more complex layout, Multi-Strategy Counter, GOAT leads to markedly higher cooperation performance with real humans, with a 38% improvement over the previous state of the art. We encourage readers to test these differences for themselves by visiting our webpage, which provides an interactive demo in which you can play Overcooked with both GOAT and the baselines https://sites.google.com/view/goat-2025/home.
(a) GOAT exploring the latent space
(b) Regret vs Minimax Exploration
Figure 5: $a)$ GOAT is exploring the latent space over training episodes, compared to red standard normal distribution cluster of GAMMA. $b)$ Comparison of Regret (Blue) with Minimax (Red) in the latent space.
@@
RQ4: Will adversarial training create a curriculum of increasingly challenging strategies by discovering them in latent space? The regret-driven adversarial framework is designed to create a dynamic curriculum where the Cooperator must continually adapt to novel challenges. GOAT identifies underperforming coordination strategies and generates increasingly complex cooperation challenges, compelling the Cooperator to explore and master new behaviors. This iterative process prevents the Cooperator from stagnating into rigid conventions, a common pitfall in self-play or population-based methods. 5(a) assesses how GOAT explores the latent space of the generative model over the whole training interval. Each point is a latent vector, and the colorbar represents the training iteration number. The 5(a) compares GOAT embeddings to the distribution of embeddings of the original GAMMA method (Liang et al., 2024) samples throughout training (red ball). We can see that at the start of training, GOAT samples latent vectors distributed throughout this ball (shown in purple). As training progresses, it chooses a particular region to exploit. After a few iterations, it moves to a new region, presumably because the Cooperator adapts to that strategy, and it is no longer able to maximize regret. Finally, at the end of the training step in yellow, GOAT has traveled across the latent space to explore particular regions (those with high regret) to find valid strategies that exploit weaknesses in the Cooperator.
@@
RQ5: Ablation study of Minimax vs. Regret. In Section 4, we hypothesized that even with a generative model in the loop, the purely adversarial minimax objective (Eq. 3) is misaligned with effectively training agents to cooperate. While the generative model can prevent sabotage behavior, because minimax focuses on worst-case scenarios, we hypothesized that a minimax adversary would still focus too much on poorly performing or incompetent partners. In this experiment, we directly compare training the GOAT adversary with either minimax (Eq. 3) or regret (Eq. 4) on the Multi-Strategy Counter Overcooked layout. We further investigate the embeddings chosen by both types of adversary throughout the training period in 5(b). We find that the minimax adversary mostly fixates on a single region (red cluster in the plot) where the policies result in low-reward behaviors or unproductive behaviors. One such example, shown as a video on our website, is where the minimax agent avoids working with its partner and converges to doing just enough not to sabotage the task. The team still receives rewards as the partner policy keeps completing the task uninterrupted. In contrast, we see that the regret objective causes GOAT to move to multiple modes in the embedding space over the course of training, exploring the latent space more effectively and thus covering a wider range of interesting behaviors. As shown in game videos on our website, it starts by doing one task, then shifts to another task, and then does both tasks simultaneously. We can also see robust training instances like intentionally coming in the way or taking roles.